OpenStudy (ashy98):

The average annual salary of the employees of a company in the year 2005 was $90,000. It increased by the same factor each year and in 2006, the average annual salary was $91,200. Let f(x) represent the average annual salary, in thousand dollars, after x years since 2005. Which of the following best represents the relationship between x and f(x)?

OpenStudy (ashy98):

A.) f(x) = 91.2(1.013)^x B.) f(x) = 90(1.013)^x C.) f(x) = 91.2(2.2)^x D.) f(x) = 90(2.2)^x